# Towards Fixing Panic Bugs for Real-world Rust Programs

### v1

@@ 3. Check whether start $>$ end if index is range:
Mutate Method Invocation
@@
3.2. Fix Pattern Mining
To comprehensively identify the causes and fix patterns for panic bugs in Rust programs, we analyze and infer data from the Rust implementation code (rustlang24:online, ). In the Rust compiler source code, panic-related error messages are encapsulated within macros such as panic_const!(). These macros and their accompanying messages provide the diverse types of panics and their fundamental causes. We compile a summary of these categories and their descriptions in Table  3. Following this categorization, we examine specific scenarios within each category. Rust compiler developers mark each panic occurrence with the annotation # Panics, which details the causes and circumstances of the panic. These annotations are often paired with # Safety or # Examples, providing either fix strategies or bug-triggering examples. This annotated information allows us to delineate code-level fix patterns for each panic type.
For example, Listing 3.2 showcases a Rust source code snippet annotated to indicate a panic bug related to the ownership mechanism. The annotation includes an example where the panic is triggered by a double borrowing of a RefCell parameter, which is a Rust smart pointer facilitating internal mutability. To prevent such panics, the annotation suggests using the try_borrow() method as an alternative approach. By systematically reviewing all documented panic instances and their proposed resolutions in the Rust source code, we collect a comprehensive list of fix strategies.
{listing}
[h] {minted}[linenos=false, fontsize=]rust /// # Panics: Panics if the value is currently /// mutably borrowed. For a non-panicking variant, /// use [‘try_borrow‘] (#method.try_borrow).
/// # Examples: An example of panic: /// let c = RefCell::new(5); /// let m = c.borrow_mut(); /// let b = c.borrow(); // this causes a panic pub fn borrow(&self) -¿ Ref¡’_, T¿ … The implementation of the borrow() function in RefCell, including documentation about potential panics.
To further delineate fix patterns, we convert code modifications into Abstract Syntax Trees (ASTs). By omitting irrelevant tokens, such as variable names, we identify similar AST transformation structures that reveal potential fixes for each i dentified root cause. In total, we collected 19 fix patterns, of which a representative categorization and corresponding examples are shown in Table 2, and the complete list of patterns can be found on our website.
Table 3. Root causes derived from panic bugs in real world.
Root Causes Code Examples
Unwrap on None/Invalid value x.unwrap() // when x = None
Mixed borrowing
x.borrow() x.borrow_mut()
Async functions wrong resume
// wrong state change
Arithmetic overflow
i32::MAX + 1
Index out of bounds
array[len + 1..]
Invalid UTF-8 boundary
&s[1..] // when s = "™";
Division/Modulus by zero
a / b // when b = 0
Assertion failed
assert!(1 == 0)
Unreachable code
unreachable!()
Others
panic!("A panic!")
3.3. Automated Panic Bugs Fixing Tool
Based on the mined fix patterns, we design a pattern-based automated panic bugs fixing tool, namely PanicKiller. PanicKiller first locates suspicious expressions based on the backtrace information output by the compiler and the dependency flow analysis of the original program. Then, combined with the semantic information of the errors, PanicKiller generates a series of patches for each fault location. Finally, PanicKiller sorts all the patches with verification and matching scores calculation, and outputs the top-5 ranked patches along with the corresponding natural language interpretation.
3.3.1. Fault Localization
When a panic occurs, Rust initiates the unwinding process, meticulously tracing back up the stack to ensure the cleanup process. The error message is further supplemented with backtrace information, and based on it, PanicKiller initially extracts suspicious locations, including file paths, column and row numbers. Considering that the precise locations of bugs might not be explicitly disclosed in the backtrace (schroter2010stacktrace, ), prior research suggests that the actual buggy location is apt to exhibit structural resemblances to the expressions detailed in the backtrace. This indicates that the exact location could be involved in other expressions and is dependent on those specified within the backtrace.
@@ 3.3.1. Fault Localization
(1) $\small Con_{i}=1-\frac{\min(Dist(\text{backtrace},e_{i}),\lambda)}{\lambda},\quad\lambda\geq 1\$
where $Dist(\text{backtrace},e_{i})$ denotes the shortest distance on the dependency graph between element $e_{i}$ and any of the elements in the backtrace. A constant $\lambda$ is used to normalize the confidence score thus ensuring it is scaled within a meaningful range. In our work, we set $\lambda=2$ according to an existing study (laura2014dep, ), which has proved that it has the optimal performance. It involves incorporating both the precise locations and the code elements dependent on those identified, ensuring a more comprehensive analysis.
In addition to the confidence score, the presence of suspicious files within the backtrace can also impact fault localization. On the one hand, files that frequently appear in the backtrace tend to be more critical to the execution path leading to the error, making them more likely to be associated with the fault. On the other hand, within an expanded backtrace, a shallower depth suggests a stronger connection to the error’s origin (wong2014stacktrace, ). Thus, PanicKiller computes the suspicion score $Loc_{i}$ for each location as follows:
@@
where $N$ denotes the number of times a particular file of $e_{i}$ occurs within the backtrace, and $D$ represents the depth of the location in the backtrace sequence. For each buggy source code and the corresponding backtrace, PanicKiller calculates the score for each element $e_{i}$ after constructing a dependency graph to determine the ranking of a suspect location.
3.3.2. Pattern-based Patch Generation
Based on the suspicious locations, PanicKiller iteratively attempts to match each location with the mined fix patterns, which are illustrated in Section 3.2. Considering that there may be multiple suspicious locations and matched patterns, resulting in multiple combinations of fixes, we sort them based on validation.
@@ 3.3.2. Pattern-based Patch Generation
Finally, PanicKiller outputs a top-5 ranked list of mixed-form fixing recommendations. To address patch interpretability issues (huang2023survey, ), PanicKiller provides each code patch alongside its natural language explanation, as detailed in Table 2. This description ensures developers receive a clear understanding of the modifications and the patterns applied in each patch.
Figure 2. A panic bug-triggering example and its patches.
@@
A running example. Figure 2 presents a code example that triggers a panic bug caused by an index out-of-bounds error. Assuming that PanicKiller has identified lines $1$ and $4$ as suspicious locations, with confidence scores of $0.7$ and $1$, respectively, as determined by Equation 2. Subsequently, PanicKiller iteratively applies each pattern to these locations. For line $1$, PanicKiller identifies a method call of unwrap(), thus the Insert Match Unwrapper pattern is applied, and patch 1 is depicted in Figure 2 (b). For line $4$, PanicKiller initially identifies a binary expression containing a subtraction operator, and the Mutate Binary Operator pattern is employed, as illustrated as patch 2 in Figure 2 (c). Further, PanicKiller analyzes the deeper structure and identifies an index expression, for which the Insert Range Checker pattern is applied to arr[num2]. Patch 3 is depicted in Figure 2 (d).
Table 4. Prioritization results for the three patches.
Patch Fix Pattern Score ($Con_{i}$+$Loc_{i}$) Validation Rank
Patch 1 Insert Match Unwrapper 1.1 (0.7 + 0.4) Test fail 3
Patch 2 Mutate Binary Operator 1.1 (1.0 + 0.1) Test pass 2
Patch 3 Insert Range Checker 1.7 (1.0 + 0.7) Test pass 1
The prioritization of patches is shown in Table 4. Patch 3 ranks first with the highest score of $1.7$, having passed the regression tests. Patch 2 is second because, although it scored the same as patch 1, it passed the regression tests while patch 1 failed. Consequently, PanicKiller outputs this ranked patch list with their interpretations, organized by scores and test results.
4. Experiment Design
In this section, we illustrate our research questions and the baseline approaches employed in our experiments. We have implemented PanicKiller with Rust, and captured Rust’s High-level Intermediate Representation (HIR) (hir:online, ), which is a special representation used in the Rust compiler, to construct the dependency graph of suspicious locations. In addition, we implement the visiting and modification of the AST nodes based on the AST manipulation library syn (syn:online, ), to generate the fix patches. All the experiments are constructed on a Ubuntu 20.04.1 LTS laptop with 11th Gen Intel(R) Core(TM) i5-11300H @3.10GHz CPU.
4.1. Research Questions
PanicKiller is designed to automatically fix Rust panic bugs. To this end, we empirically explore the following research questions (RQ):
RQ1 (Fault Localization): To what extent does PanicKiller provide reliable fault localization results for panic bugs fixing? In this RQ, we aim to evaluate the accuracy of our bug localization approach based on backtrace information and dependency analysis. We categorize the real data in Panic4R into small-scale datasets ($<20$K LoC) and large-scale datasets ($\geq 20$K LoC) based on the size of the source code. We present the accuracy of various localization approaches at top-1, top-3, and top-5 levels. The accuracy is calculated based on three granularity levels: files, statements, and expressions.
RQ2 (Fixing Effectiveness): How effective is PanicKiller for fixing panic bugs compared to state-of-the-art LLM-based techniques? To evaluate PanicKiller’s effectiveness, we first verified the accuracy of patches it generated on Panic4R. Additionally, we conducted comparative experiments to assess patch generation accuracy when correct error localization information was given to both PanicKiller and baseline methods. Referring to existing research on APR (qi2015plausible, ), we separately verify the proportion of fixes that are plausible, i.e., the panic bug is successfully fixed, but some of the regression test cases fail, and the proportion of correct fixes, i.e., the panic bug is successfully fixed, and the regression test cases all pass.
RQ3 (Efficiency): What is the fixing efficiency of PanicKiller compared with traditional manual repair methods? In this RQ, we measure the time required to apply PanicKiller for fixing each target, and compare it with the duration from when the issue was reported to the PR was merged. This comparison aims to assess the extent to which our proposed approach enhances the efficiency of bug repairing.
Table 5. The correctness of tools on real and synthetic datasets, respectively, with different granularities of error localization.
Tool Panic4R-Small (61) Panic4R-Large (41) Total (102)
File Stmt Expr File Stmt Expr File Stmt Expr
RQ4 (Case Study): Can PanicKiller be applied to the real-world Rust projects and fix practical panic bugs? To evaluate the practical effectiveness, we apply PanicKiller to fix real-world issues in GitHub, which are not included in Panic4R. For open issues, we submit a PR with the fix patch generated by PanicKiller, and we track whether they are confirmed and merged by developers. For closed issues, we use TF-IDF (sparck1972statistical, ) to calculate the similarity between the generated patch with the official patch. If the score exceeds $0.75$, and the panic is fixed with the regression test passed, we consider they are similar. If the PR is accepted by project developers or if our generated patch is similar to the official patch, it could further demonstrate that PanicKiller can be employed to real-world application scenarios and provide fixing guidance for large-scale real-world projects.
4.2. Baseline Approaches
In bug localization, while existing techniques typically rely on additional natural language descriptions from bug reports or test suites (yoo2012regression, ), our tool, PanicKiller, solely utilizes the error message from the Rust compiler and the accompanying backtrace. The absence of extra data makes traditional techniques unsuitable as baselines for our approach. Furthermore, since these techniques are designed for languages like C, C++, Java, and Python, they are not tailored to handle panic errors in Rust, which have distinct error message characteristics. Consequently, we devised three baseline methods based on conventional localization approaches: (1) panic: Directly using the location indicated in the panic error message. (2) random: Randomly selecting a suspicious location from the backtrace. (3) similarity: Calculating textual similarity between the error message and each code location; the highest similarity score indicates the fault location.
@@ 4.2. Baseline Approaches
To our knowledge, no existing tools specifically address bug localization or the automated repair of Rust panic bugs. We therefore benchmark against the state-of-the-art LLM-based technique, ChatGPT. Due to ChatGPT-3.5’s limitations in handling files, we utilize the advanced and commercial ChatGPT-4.0 as our primary baseline. For each target, we upload a zip archive containing the entire Rust project, the test case triggering the panic, and relevant panic information. We then use a consistent prompt template, as shown below:
@@
5. Result Analysis
In this section, we present the experiment results on automatically fixing panic bugs and then analyze the effectiveness and efficiency of our proposed tool PanicKiller.
5.1. Answer to RQ1: Fault Localization
@@
To comprehensively assess the impact of fault localization through various techniques, we employ three granularity levels for locating panics: file, statement, and expression levels. As illustrated in Table 5, the localization accuracy of PanicKiller at all three granularities is higher than the other baselines, which indicates the effectiveness of PanicKiller. In addition, all methods perform better on the small-scale dataset compared with the large-scale dataset. This result is as expected, because on large-scale projects, the stack unfolding information could be more complex and harder to analyze the dependency for fault localization.
From the results, the localization accuracy of ChatGPT4 is basically the same for single and multi-round conversations due to a limited understanding of an entire Rust project. Throughout our experiments, we have encountered ChatGPT failing to parse the file and only outputting some broad suggestions, which is not useful for practical fault localization. Interestingly, the effectiveness of the random method is relatively close to PanicKiller, especially on small-scale datasets and top-5 prediction results, which suggests that if the localization is in the stack-expanded list, random selection can be effective, but as soon as in-depth dependency analyses are required +on large-scale dataset?, the random method fails outright. The similarity-based selection techniques are the least effective in terms of fine-grained localization accuracy, implying that relying on natural language alone does not accurately capture the semantic information of complex practical code, resulting in low localization accuracy.
5.2. Answer to RQ2: Fixing Effectiveness
Because there is no existing APR tool for Rust panic bugs, we focus on comparison the fixing effectiveness with ChatGPT-4. The evaluation results are shown in Table 6. Overall, we can conclude that PanicKiller significantly surpasses the performance of automated repairs conducted by GPT-4. Notably, within the large-scale dataset, GPT-4 fails to generate any viable patches, the reason of which is closely tied to GPT-4’s challenges with fault localization, as elaborated in Section 5.1. Moreover, we find that LLM-based methods frequently overlooks the code context when generating patches, resulting in the generated program failing to compile instead. For example, it may attempt to resolve panics by altering a method’s return type directly or by incorporating new logic, which results in both compilation and semantic errors.
Table 6. The number of correct patches and correctness of different tools on Panic4R, respectively.
@@ 5.2. Answer to RQ2: Fixing Effectiveness
Plausible Correct Plausible Correct Plausible Correct
Top-1 GPT-4 6 9.8% 6 9.8% 0 0% 0 0% 6 5.9% 6 5.9%
GPT4-multi 8 13.1% 8 13.1% 0 0% 0 0% 8 7.8% 8 7.8%
PanicKiller 29 47.5% 19 31.1% 10 24.4% 7 17.1% 39 38.2% 26 25.5%
Top-3 GPT-4 6 9.8% 6 9.8% 0 0% 0 0% 6 5.9% 6 5.9%
GPT4-multi 8 13.1% 8 13.1% 0 0% 0 0% 8 7.8% 8 7.8%
PanicKiller 32 52.5% 23 37.7% 13 31.7% 10 24.4% 45 41.1% 33 32.4%
Top-5 GPT-4 7 11.5% 7 11.5% 0 0% 0 0% 7 6.9% 7 6.9%
GPT4-multi 9 14.8% 9 14.8% 0 0% 0 0% 9 8.8% 9 8.8%
PanicKiller 35 57.4% 25 41.0% 14 34.1% 10 24.4% 49 48.0% 35 34.3%

### v2

@@ 3. Check whether start $>$ end if index is range:
Mutate Method
Invocation
@@
5. Automated Panic Bugs Fixing Tool
Based on the mined fix patterns, we design a pattern-based automated panic bugs fixing tool, namely PanicKiller. PanicKiller first locates suspicious expressions based on the stacktrace information output by the compiler and the dependency flow analysis of the original program. Then, combined with the semantic information of the errors, PanicKiller generates a series of patches for each fault location. Finally, PanicKiller sorts all the patches with verification and matching scores calculation, and outputs the top-5 ranked patches along with the corresponding natural language interpretation.
5.1. Fault Localization
When a panic occurs, Rust initiates the unwinding process, meticulously tracing back up the stack to ensure the cleanup process. The error message is further supplemented with stacktrace information, and based on it, PanicKiller initially extracts suspicious locations, including file paths, column and row numbers. Considering that the precise locations of bugs might not be explicitly disclosed in the stacktrace (Schroter et al., 2010), prior research suggests that the actual buggy location is apt to exhibit structural resemblances to the expressions detailed in the stacktrace. This indicates that the exact location could be involved in other expressions and is dependent on those specified within the stacktrace.
@@ 5.1. Fault Localization
(1) $\small Con_{i}=1-\frac{\min(Dist(\text{stacktrace},e_{i}),\lambda)}{\lambda},\quad\lambda\geq 1\$
where $Dist(\text{stacktrace},e_{i})$ denotes the shortest distance on the dependency graph between element $e_{i}$ and any of the elements in the stacktrace. A constant $\lambda$ is used to normalize the confidence score thus ensuring it is scaled within a meaningful range. In our work, we set $\lambda=2$ according to an existing study (Moreno et al., 2014), which has proved that it has the optimal performance. It involves incorporating both the precise locations and the code elements dependent on those identified, ensuring a more comprehensive analysis.
In addition to the confidence score, the presence of suspicious files within the stacktrace can also impact fault localization. On the one hand, files that frequently appear in the stacktrace tend to be more critical to the execution path leading to the error, making them more likely to be associated with the fault. On the other hand, within an expanded stacktrace, a shallower depth suggests a stronger connection to the error’s origin (Wong et al., 2014). Thus, PanicKiller computes the suspicion score $Loc_{i}$ for each location as follows:
@@
where $N$ denotes the number of times a particular file of $e_{i}$ occurs within the stacktrace, and $D$ represents the depth of the location in the stacktrace sequence. For each buggy source code and the corresponding stacktrace, PanicKiller calculates the score for each element $e_{i}$ after constructing a dependency graph to determine the ranking of a suspect location.
5.2. Pattern-based Patch Generation
Based on the suspicious locations, PanicKiller iteratively attempts to match each location with the mined fix patterns, which are illustrated in Section 4. Considering that there may be multiple suspicious locations and matched patterns, resulting in multiple combinations of fixes, we sort them based on validation.
@@ 5.2. Pattern-based Patch Generation
Finally, PanicKiller outputs a top-5 ranked list of mixed-form fixing recommendations. To address patch interpretability issues (Huang et al., 2023), PanicKiller provides each code patch alongside its natural language explanation, as detailed in Table 3. This description ensures developers receive a clear understanding of the modifications and the patterns applied in each patch.
Figure 5. A panic bug-triggering example and its corresponding patches.
@@
A running example. Figure 5 (a) presents a code example that triggers a panic bug caused by an index out-of-bounds error. Assuming that PanicKiller has identified lines $1$ and $4$ as suspicious locations, with confidence scores of $0.7$ and $1$, respectively, as determined by Equation 2. Subsequently, PanicKiller iteratively applies each pattern to these locations. For line $1$, PanicKiller identifies a method call of unwrap(), thus the Insert Match Unwrapper pattern is applied, and patch 1 is depicted in Figure 5 (b). For line $4$, PanicKiller initially identifies a binary expression containing a subtraction operator, and the Mutate Binary Operator pattern is employed, as illustrated as patch 2 in Figure 5 (c). Further, PanicKiller analyzes the deeper structure and identifies an index expression, for which the Insert Range Checker pattern is applied to arr[num2]. Patch 3 is depicted in Figure 5 (d). The prioritization of patches is shown in Figure 5 (e). Patch 3 ranks first with the highest score of $1.7$, having passed the regression tests. Patch 2 is second because, although it scored the same as patch 1, it passed the regression tests while patch 1 failed. Consequently, PanicKiller outputs this ranked patch list with their interpretations, organized by scores and test results.
5.3. Tool Evaluation
@@
To the best of our knowledge, there currently exists no dedicated tool for the automated repair of Rust panic bugs. Therefore, we conducted an evaluation on the Panic4R dataset to validate the effectiveness of PanicKiller. We mainly compare the effectiveness of PanicKiller with LLM-based approaches. We follow a standardized process when applying ChatGPT 4.0 as a baseline method for fault localization and program repair. For each target, we upload a zip archive containing the entire Rust project, the test case triggering the panic, and relevant panic information. We then use a consistent prompt template, as shown below:
@@
5.3.1. Effectiveness of Fault Localization
For the task of bug localization, while existing techniques typically rely on additional natural language descriptions from bug reports or test suites (Yoo and Harman, 2012), our tool, PanicKiller, solely utilizes the error message from the Rust compiler and the accompanying stacktrace information. The absence of extra data makes some existing techniques unsuitable as baselines for our approach. Consequently, we select three categories of general methods as baseline techniques to compare with the fault localization capabilities of PanicKiller.
(1) Conventional Localization Approaches. When only error messages and stacktraces are available, fault localization can be deduced from these elements(Jiang et al., 2012; Moreno et al., 2014; Wong et al., 2014).
  • Panic: When a panic occurs, the compiler outputs an error message indicating the specific line of code where the exception occurred. This line is then considered as the fault location.
  • Random: The compiler outputs stacktrace information that includes a list of suspicious code locations. From this list, a location is randomly selected as the detected fault location.
  • Similarity: We evaluate the textual error message and identify the fault location by finding the code line with the highest similarity score.
(2) Spectrum-based Localization Approaches. Spectrum-based fault localization (SBFL) (Abreu et al., 2007) approaches analyze the execution traces of a program by examining which components of the program were executed when a failure occurred and which were not. By applying statistical analysis to these execution spectra, SBFL effectively pinpoints the components that are most likely to be responsible for software failures. In our experiments, we collect test cases for each crate in Panic4R, the majority of which passed; those that triggered a panic bug were classified as failed cases. Based on the calculated scores from the spectral analysis, we sort all the lines of code and then select the top 5 locations as the fault localization results.
  • Tarantula: The Tarantula method (Jones et al., 2002) calculates the suspicious location by comparing the ratios of passed to failed test cases that execute a given code line. A higher ratio of failed to total test executions in a component increases its likelihood of being faulty. We rank the ratio of each code line as the most likely fault candidates.
  • Ochiai: The Ochiai algorithm (Meyer et al., 2004) applies the cosine similarity measure between failed test cases and the execution of given code lines. The score is derived from the intersection of failing test cases with the components they execute, normalized by the square root of the product of the total number of failures and the number of times a component is executed during those failures.
(3) LLM-based Localization Approaches. It is feasible to directly submit source code, error messages, and requirements to an LLM.
  • ChatGPT-4 (GPT-4): Due to ChatGPT-3.5’s limitations in handling files, we apply the most popular commercial software ChatGPT 4.0 as a baseline.
  • Multiple rounds of inquiries with ChatGPT-4 (GPT4-multi): When interacting with ChatGPT 4.0, we inquire multiple rounds with additional questions manually.
@@ 5.3.1. Effectiveness of Fault Localization
To comprehensively assess the effectiveness of fault localization through various techniques, we employ three granularity levels for locating panics: file, statement, and expression levels. As illustrated in Table 4, the localization accuracy of PanicKiller at all three granularities is higher than the other baselines, which indicates the effectiveness of PanicKiller. In addition, all methods perform better on the small-scale dataset compared with the large-scale dataset. This result is as expected, because on large-scale projects, the stack unfolding information could be more complex and harder to analyze the dependency for fault localization. The effectiveness of the random method is relatively close to PanicKiller, especially on small-scale datasets and top-5 prediction results, which suggests that if the localization is in the stack-expanded list, random selection can be effective, but as soon as in-depth dependency analyses are required on large-scale dataset, the random method fails outright. The similarity-based selection techniques are the least effective in terms of fine-grained localization accuracy, implying that relying on natural language alone does not accurately capture the semantic information of complex practical code, resulting in low localization accuracy.
As for comparing with LLM-based approaches, the localization accuracy of ChatGPT4 is basically the same for single and multi-round conversations due to a limited understanding of an entire Rust project. This limitation is evident in the significant drop in success rate as the size of the Rust project increases. Throughout our experiments, we observed instances where ChatGPT failed to parse files accurately, providing only general suggestions that are impractical for effective fault localization. These observations highlight the inconsistency and limitations of relying on LLMs for accurate bug localization in complex codebases.
For spectrum-based methods, both Ochiai and Tarantula methods exhibited low accuracy, particularly within larger datasets, where Ochiai, for instance, showed zero accuracy in the top-1 ranking across all levels. This may because that panic bugs in Rust program constitute a unique category of errors, usually not originating from traditional issues in loops or conditional logic. The unit test cases within each crate aim to validate logical functionality, thus testing pass and fail conditions that may not align with detecting panic-related bugs. This mismatch could reduce the effectiveness of spectrum-based techniques. In addition, facing with real-world and large-scale programs that contain thousands of lines code, applying spectrum-based techniques incurs considerable time overheads, further limiting their practicality. For example, calculating the spectrum score for a project could takes more than 3 hours in average for the large-scale dataset, while PanicKiller takes about 1 minute in average.
Table 4. The correctness of tools on Panic4R datasets, with different granularities of error localization.
@@
File Stmt Expr File Stmt Expr File Stmt Expr
Ochiai 0 (0%) 0 (0%) 0 (0%) 0 (0%) 0 (0%) 0 (0%) 0 (0%) 0 (0%) 0 (0%)
Tarantula 1 (1.6%) 1 (1.6%) 1 (1.6%) 0 (0%) 0 (0%) 0 (0%) 1 (1.0%) 1 (1.0%) 1 (1.0%)
Ochiai 3 (5.0%) 3 (5.0%) 3 (5.0%) 0 (0%) 0 (0%) 0 (0%) 3 (2.9%) 3 (2.9%) 3 (2.9%)
Tarantula 2 (3.3%) 2 (3.3%) 2 (3.3%) 0 (0%) 0 (0%) 0 (0%) 2 (2.0%) 2 (2.0%) 2 (2.0%)
Ochiai 9 (14.8%) 9 (14.8%) 9 (14.8%) 0 (0%) 0 (0%) 0 (0%) 9 (8.8%) 9 (8.8%) 9 (8.8%)
Tarantula 6 (9.8%) 6 (9.8%) 6 (9.8%) 1 (2.4%) 1 (2.4%) 1 (2.4%) 7 (6.9%) 7 (6.9%) 7 (6.9%)
5.3.2. Fixing Effectiveness
Referring to existing research on APR (Qi et al., 2015), we categorize the generated patches into three types and verify their proportion: (1) panic-eliminated patches, where the panic bug is resolved but some regression test cases fail; (2) plausible patches, where the panic bug is successfully fixed and all test cases pass; and (3) correct patches, which are plausible patches that have also been manually verified for semantic correctness. The evaluation results are shown in Table 5. Overall, we can conclude that PanicKiller significantly surpasses the performance of automated repairs conducted by GPT-4. Notably, within the large-scale dataset, GPT-4 fails to generate any viable patches, the reason of which is closely tied to GPT-4’s challenges with fault localization, as elaborated in Section 5.3.1. Moreover, we find that LLM-based methods frequently overlooks the code context when generating patches, resulting in the generated program failing to compile instead. For example, it may attempt to resolve panics by altering a method’s return type directly or by incorporating new logic, which results in both compilation and semantic errors.
Table 5. The number of panic-eliminated/plausible/correct patches of different tools on Panic4R.
Tool Panic4R-Small (61) Panic4R-Large (41) Total (102)
Eliminated Plausible Correct Eliminated Plausible Correct Eliminated Plausible Correct
Top-1 GPT-4 6 9.8% 6 9.8% 0 0% 0 0% 0 0% 0 0% 6 5.9% 6 5.9% 0 0%
GPT4-multi 8 13.1% 8 13.1% 0 0% 0 0% 0 0% 0 0% 8 7.8% 8 7.8% 0 0%
PanicKiller 29 47.5% 19 31.1% 4 6.6% 10 24.4% 7 17.1% 4 9.8% 39 38.2% 26 25.5% 8 7.8%
Top-3 GPT-4 6 9.8% 6 9.8% 1 1.6% 0 0% 0 0% 0 0% 6 5.9% 6 5.9% 1 1.0%
GPT4-multi 8 13.1% 8 13.1% 1 1.6% 0 0% 0 0% 0 0% 8 7.8% 8 7.8% 1 1.0%
PanicKiller 32 52.5% 23 37.7% 12 19.7% 13 31.7% 10 24.4% 7 17.1% 45 41.1% 33 32.4% 19 18.6%
Top-5 GPT-4 7 11.5% 7 11.5% 1 1.6% 0 0% 0 0% 0 0% 7 6.9% 7 6.9% 1 1.0%
GPT4-multi 9 14.8% 9 14.8% 1 1.6% 0 0% 0 0% 0 0% 9 8.8% 9 8.8% 1 1.0%
PanicKiller 35 57.4% 25 41.0% 12 19.7% 14 34.1% 10 24.4% 7 17.1% 49 48.0% 35 34.3% 19 18.6%